### v4

# Exploring the presence of a fifth force at the Galactic Center

Abstract
Aims. We investigate the presence of a Yukawa-like correction to Newtonian gravity at the Galactic Center, leading to a new upper limit on the intensity of such a correction.
Methods. We performed a Markov chain Monte Carlo (MCMC) analysis using the astrometric and spectroscopic data of star S22 collected at the Very Large Telescope by GRAVITY, NACO, and SINFONI instruments, covering the period from $1992$ to $2022$.
Results. The precision of the GRAVITY instrument allows us to derive the most stringent upper limit at the Galactic Center for the intensity of the Yukawa contribution ($\propto\,\alpha e^{-\lambda r}$) of $|\alpha|<0.003$ for a scale length of $\lambda=3\cdot 10^{13}\,\rm m\,(\sim 200\,\rm AU)$. This is an improvement on all estimates obtained in previous works by roughly one order of magnitude.
Key Words.:
black holes physics – Galaxy:centre – gravitation
1 Introduction
General relativity (GR) is the most widely recognized theory of gravity today. Its predictions have been extensively tested on Solar System scales and using gravitational waves emission by black holes (BHs) and binary pulsars (Will, 2014, 2018b; Nitz et al., 2021). Until now, no significant deviation from GR has been detected in any of these observations. However, it is also known that beyond the regime we are currently able to test experimentally, GR is generally considered insufficient.
Cosmological observations indicate an expanding Universe whose acceleration can only be explained by introducing a cosmological constant ad hoc, which lacks a solid theoretical explanation and raises several issues (Weinberg, 1989; Peebles and Ratra, 2003). Other observational evidence, such as the rotational curve of galaxies (van Albada et al., 1985; Salucci, 2019) or gravitational lensing effects (Massey et al., 2010) indicate the presence of a dark massive component of the Universe, whose nature is still unknown. Furthermore, it is well known that GR lacks a quantum description at high-energy scales and several attempts have been made to create a theory valid at all scales (for a review on the state-of-the-art approaches, see, e.g., Esposito (2011); Kiefer (2023)).
One way to address these inconsistencies between theory and experiments is to directly modify GR, giving rise to a plethora of possible extended theories of gravity (ETG). In particular, a Yukawa-like interaction emerges quite naturally in the weak field limit of several ETGs; for instance, scalar-tensor-vector theories (Moffat, 2006), massive gravity theories (Visser, 1998; Hinterbichler, 2012), theories in higher dimensions with Kaluza-Klein compactification (Bars and Visser, 1986; Hoyle et al., 2001), massive Brans-Dicke theories (Perivolaropoulos, 2010; Alsing et al., 2012), or $f(R)$ theories (Capozziello et al., 2015). However, the so-called ”fifth-force” scenario has also made an appearance in a number of specific dark matter models (Frieman and Gradwohl, 1991; Gradwohl and Frieman, 1992; Carroll et al., 2009).
Due to the importance that a modification of Newtonian gravity would have on our understanding of the Universe, the presence of a Yukawa-like contribution has been repeatedly investigated in the past. The fifth-force intensity is well constrained at Solar System scales via the motion of planets (Konopliv et al., 2011; Hees et al., 2014; Bergé, 2017; Will, 2018a; Shankaranarayanan and Johnson, 2022), from the Lunar Laser Ranging experiment (Hofmann and Müller, 2018) and also by making use of the planetary ephemerides (Mariani et al., 2023; Fienga and Minazzoli, 2024). Recent constraints have been obtained from asteroids tracking (Tsai et al., 2023, 2024) and test of the weak equivalence principle (Touboul and others, 2022).
The discovery of orbiting stars around the Galactic Center (GC) (Eckart and Genzel, 1996; Schödel et al., 2002; Ghez et al., 2003; Gillessen et al., 2009b, a; Sabha and others, 2012), all located within one arcsecond distance from the supermassive black hole (SMBH) Sagittarius A (Sgr A), allows us to test GR in a completely different environment from the Solar System.
The importance of looking for a fifth force in the GC lies in the fact that many ETGs that predict a Yukawa-like term also display a screening mechanism that suppresses the fifth force contribution at Solar System scales and prevents its detection. This would explain why it would be yet unobserved, while its effect may be different around SMBHs.
The current constraints on the intensity of a fifth force in the GC come from the analysis of SgrA’s shadow by the Event Horizon Telescope (Vagnozzi and others, 2023), from the measurement of the Schwarzschild precession in S22 motion (GRAVITY Collaboration, 2020; Jovanović et al., 2023, 2024b, 2024a) and from the analysis of S-stars publicly available (or mock) data (Borka et al., 2013; Capozziello et al., 2014; Borka et al., 2021; Zakharov et al., 2016, 2018; de Martino et al., 2021; Della Monica et al., 2022), while also including the presence of a (expected) bulk mass distribution around Sgr A (Jovanović et al., 2021). For a complete and comprehensive review on astrophysical and theoretical constraints at the GC, we refer to de Laurentis et al. (2023).
In the context of S-stars, a previous work by Hees and others (2017), with a full analysis of the S22 data, showed that the intensity of such a contribution cannot exceed $\alpha\sim 0.01$ at scales comparable to the S22-SgrA distance.
In this Letter, we use the astrometric and spectroscopic measurements of the star S22 collected at the Very Large Telescope (VLT) by GRAVITY, NACO, and SINFONI to constrain the intensity of a possible Yukawa correction at the GC. Although we do not expect our results to consistently deviate from the estimates obtained in the aforementioned literature, a complete analysis of S22 including GRAVITY data, which dominate the $\chi^{2}$ due to their very small uncertainties, is still lacking. As we show here, the precision of the GRAVITY instrument allows us to place a significantly stronger constraint than the previous estimates.
2 Observations
The set of available data, $D,$ can be organized according to the following criteria:
  • a)
    Astrometric data $\rm DEC$, $\rm R.A.$
    • 128 data points collected using both the SHARP camera at the New Technology Telescope between 1992 and 2002 ($\sim$ 10 data points, accuracy of $\approx 4\,\rm mas$) and the NACO imager at the VLT between 2002 and 2019 (118 data points, accuracy of $\approx 0.5\,\rm mas$);
    • 76 data points collected by GRAVITY at the VLT between 2016 and April 2022 (accuracy of $\approx 50\,\rm\mu as$).
  • b)
    Spectroscopic data $V_{R}$
    • 102 data points collected by SINFONI at the VLT (100 points) and NIRC2 at Keck (2 points) collected between 2000 and March 2022 (accuracy in good conditions of $\approx 10-15\,\rm km/s$).
3 Yukawa correction to Newtonian force
3.1 Model
The potential we aim to test takes the following form:
$U=-\frac{GM}{r}\left(1+|\alpha|e^{-r/\lambda}\right)\,$, (1)
where $\alpha$ represents the strength of interaction and $\lambda$ is a scale parameter, which depends on the specific theory considered. For example, when new massive fields are included in the theory, $\lambda$ represents the Compton wavelength of the field, which is related to the mass by $m_{\varphi}=h/c\lambda$, where $h$ is the Planck constant.
In comparison to previous works, here the Schwarzschild precession in the S22 orbit is also included, since it has been formally detected at a $10\sigma$ confidence level by the GRAVITY Collaboration (GRAVITY Collaboration, 2020, 2024).
Although a formal parametrized post-Newtonian (PN) treatment is not possible when a massive field is included in the action (Alsing et al., 2012; Poisson and Will, 2012), we can still derive the equations of motion of a test particle, assuming the parametrized PN parameters to be $\gamma=\beta=1$. This latter assumption is valid for ETGs that are indistinguishable from GR at 11PN order and it is supported by different experimental observations, including those carried out for the GC (see, e.g., Will (2018b); Hofmann and Müller (2018); GRAVITY Collaboration (2020); Iorio (2024)). On the other hand, non-metric theories of gravity or specific subclasses of f(R) and scalar-tensor theories, with parametrized PN parameters that significantly stray from unity, are excluded from this analysis (Will, 2014).
The total acceleration experienced by the star is
$\mathbf{a}_{\rm TOT}=\mathbf{a}_{\rm New}+\mathbf{a}_{\rm Yuk}+\mathbf{a}_{\rm 1PN}\,$, (2)
where $\mathbf{a}_{\rm New}+\mathbf{a}_{\rm Yuk}$ are derived from the potential in Eq. (1) and
$\boldsymbol{a}_{1\rm PN}=\frac{GM}{c^{2}r^{2}}\left[\left(\frac{4GM}{r}-v^{2}\right)\frac{\boldsymbol{r}}{r}+4\dot{r}\boldsymbol{v}\right]\,$, (3)
[…]
The above expression coincides with the 11PN acceleration derived in Alves et al. (2024) for the two-body problem in massive Brans-Dicke theory. Section 3.4 will be devoted to a comparison of Eq. (3) with results developed in the literature when extra massive degrees of freedom are included in the theory.
3.2 Method
The numerical integration of the equations of motion was performed using a Runge-Kutta $4(5)$ method, with further details reported in Appendix A. During the fit of the S22 data, the Yukawa length scale, $\lambda$, was kept fixed, with values set between $10^{12}\leq\lambda\leq 10^{15}\,\rm m$, while the intensity, $\alpha$, was allowed to vary together with other parameters describing the system.
Specifically, the set of parameters is given by:
$\Theta_{i}=\{e,a_{\rm sma},\Omega_{\rm orb},i_{\rm orb},\omega_{\rm orb},t_{p},R_{0},M,x_{0},y_{0},v_{x_{0}},v_{y_{0}},v_{z_{0}},\alpha\},$ (4)
where $e$ is the eccentricity and $a_{\rm sma}$ the semi major axis of the star S22, while $\Omega_{\rm orb}$, $i_{\rm orb}$, and $\omega_{\rm orb}$ are the three angles used to project the star’s orbital frame into the observer reference frame using the procedure reported in Appendix B.1. Then, $t_{p}$ is the time of pericenter passage, while $M$ and $R_{0}$ are the SMBH mass and the GC distance, respectively. The additional parameters $\{x_{0},y_{0},v_{x_{0}},v_{y_{0}},v_{z_{0}}\}$ characterize the NACO/SINFONI data reference frame with respect to Sgr A (Plewa et al., 2015).
To fit the S22 data, we perform a Markov chain Monte Carlo (MCMC) analysis using the Python package emcee (Foreman-Mackey et al., 2013). The log-likelihood is given by
$\ln\mathscr{L}=\ln\mathscr{L}_{\rm pos}+\ln\mathscr{L}_{\rm vel}\,$, (5)
where
$\ln\mathscr{L}_{\rm pos}=-\sum_{i=1}^{N}\left[\frac{(\rm DEC_{i}-\rm DEC_{\rm model,i})^{2}}{\sigma_{\rm DEC_{i}}^{2}}+\frac{(\rm R.A._{i}-\rm R.A._{\rm model,i})^{2}}{\sigma_{\rm R.A._{i}}^{2}}\right]\,$, (6)
and
$\ln\mathscr{L}_{\rm vel}=-\sum_{i=1}^{N}\frac{(V_{R,i}-V_{\rm model,i})^{2}}{\sigma_{V_{R,i}}^{2}}\,.$ (7)
The priors are listed in Tables 1 and 1. Uniform priors were used for the physical parameters, that is, we only imposed physically motivated bounds, while Gaussian priors were implemented for the offset parameters, since the latter have been well constrained by an independent previous work and were not expected to change (Plewa et al., 2015). The initial points $\Theta_{i}^{0}$ in the MCMC are chosen to be the most recent best fit parameters of S22 orbit reported in the literature (GRAVITY Collaboration, 2024).
Table 1: Uniform priors used in the MCMC analysis.
Parameter $\Theta_{i}^{0}$ Lower bound Upper bound
$e$ 0.88441 0.83 0.93
$a_{\rm sma}$ [”] 0.12497 0.119 0.132
$i_{\rm orb}\,[^{\circ}]$ $134.69241$ 100 $150$
$\omega_{\rm orb}\,[^{\circ}]$ $66.28411$ 40 $90$
$\Omega_{\rm orb}\,[^{\circ}]$ $228.19245$ $200$ $250$
$t_{p}$ [yrs] 2018.37902 2018 2019
$M\,[10^{6}\,M_{\odot}]$ $4.29950$ 4.1 4.8
$R_{0}\,\rm[10^{3}\,pc]$ 8.27795 8.1 8.9
[…]
Parameter $\Theta_{i}^{0}$ $\xi$ $\sigma$
$x_{0}\,\rm[mas]$ -0.244 -0.055 0.25
$y_{0}\,\rm[mas]$ -0.618 -0.570 0.15
$v_{x_{0}}\,\rm[mas/yr]$ 0.059 0.063 0.0066
$v_{y_{0}}\,\rm[mas/yr]$ 0.074 0.032 0.019
$v_{z_{0}}\,\rm[km/s]$ -2.455 0 5
In the sampling phase of the MCMC implementation, we used 64 walkers and $10^{5}$ iterations. The burning-in phase was skipped and the last $80\%$ of the chains was used to compute the mean and standard deviation of the posterior distributions of the parameters. The convergence of the MCMC analysis was ensured by means of the autocorrelation time $\tau_{c}$, that is, we ran $N$ iterations such that $N\gg 50\,\tau_{c}$.
3.3 Results
We can classify three different regimes in the posterior distributions $P(|\alpha||D)$, according to the value of $\lambda$ with respect to the orbital range of S22, which is $1.7\cdot 10^{13}\,\rm m\lesssim r_{s2}\lesssim 1.5\cdot 10^{14}\,\rm m$. When $\lambda\ll r_{s2}$, the acceleration is no longer dependent on the parameter $\alpha$ and no meaningful constraints can be obtained in this regime. The small difference in the $95\%$ upper limit on $|\alpha|$ with the UCLA group resides in the different model implemented to fit the data.
In particular, Hees and others (2017) considered no Schwarzschild precession but an extended mass with power law distribution instead, which is absent from our work. The reason behind this choice comes from the fact that the presence of a spherically symmetric mass distribution around SgrA has been extensively tested by the GRAVITY Collaboration (GRAVITY Collaboration, 2018, 2022, 2024), finding a stringent upper limit of $M_{\rm ext}\lesssim 1200\,M_{\odot}$. Taking into account this current upper limit, we decided to neglect the presence of an extended mass, as it would not alter our results.
When $\lambda\sim r_{s2}$, the best constraints on $|\alpha|$ can be obtained, determining the most stringent upper limit of $|\alpha|<0.003$ for $\lambda=3\cdot 10^{13}\,\rm m\sim 200\,\rm AU$. This limit improves the previous estimate of Hees and others (2017), which reported $|\alpha|<0.016$ for $\lambda=150\,\rm AU\sim 2.2\cdot 10^{13}\,\rm m$, stressing the importance of the precision of the GRAVITY instrument.
Finally, when $\lambda\gg r_{s2}$, the only component left in the equations of motion is the monopolar term $M(1+\alpha)/r$. This corresponds to a simple rescaling of the mass term and, hence, in this regime $M$ and $\alpha$ are completely degenerate and they cannot be constrained separately. To obtain the upper limit on $\alpha$ in this regime, the bounds on $M$ in Table 1 have been extended to $M\in(10^{-4},10^{4})\cdot 10^{6}\,M_{\odot}$, and the same bounds have been used for $\alpha$.
A summary of the above results is reported in Figure 2, where the $95\%$ confidence interval on $|\alpha|$ as function of $\lambda$ is shown. Those confidence intervals are estimated as three times the standard deviation when the posterior distributions are normal or by computing the upper limit that corresponds to $95\%$ of the area below the curve when the distributions have different shapes. We note that GRAVITY data produce an overall improvement of roughly one order of magnitude over the entire parameter space tested.
If we assume that the gravitational interaction is mediated by a massive boson as in massive gravity theories (where $\alpha=1$), the length scale $\lambda$ corresponds to the Compton wavelength of the particle; hence, an upper limit on the graviton’s mass can be derived. Since $\alpha=1$ is excluded at $95\%$ confidence level for $\lambda\lesssim 8\cdot 10^{14}\,\rm m$, this lower bound on the wavelength, $\lambda$, can be translated into an upper limit on the graviton mass, corresponding to $m_{g}\lesssim 2.5\cdot 10^{-22}\,\rm eV$.
In the regime where $\lambda\sim r_{s2}$, when no 11PN acceleration is included, the presence of the Yukawa term in the equation of motion induces a prograde precession comparable to the Schwarzschild one. This is shown in Figure 1 for $\lambda=10^{13}\,\rm m$, where we can see that the posterior distribution $P(\alpha|D)$ is a Gaussian with mean around $\alpha\sim 0.0026$, as opposed to the 11PN posterior.
Following Adkins and McDonnell (2007), we can compute the precession angle induced by a potential in a full orbit as
$\Delta\phi_{p}=-\frac{2L}{GMe^{2}}\int_{-1}^{1}\frac{dz\,z}{\sqrt{1-z^{2}}}\frac{dU(z)}{dz},$ (8)
where $U(z)$ is the perturbing potential evaluated at radius $r=L/(1+ez)$ with $L=a_{\rm sma}(1-e^{2})$. For $\alpha=0.0026$, this corresponds to $\Delta\phi_{p}\sim 0.13^{\circ}$. Taking the most up-to-date value of the Schwarzschild precession reported in GRAVITY Collaboration (2024), $\Delta\phi_{\rm Sch}=12.1^{\prime}\times(0.911\pm 0.13)=(0.18\pm 0.03)^{\circ}$, we can see that the precession angle induced by the Yukawa potential is compatible, within the $2\sigma$ uncertainties, with $\Delta\phi_{\rm Sch}$.
Figure 1: Comparison of the posterior distributions $P(|\alpha||D)$ between the Keplerian model (black curve) and the 11PN model (red curve).
Figure 2: $95\%$ confidence level on $|\alpha|$ obtained in this work (red dots), compared with previous estimates by Hees and others (2017) (blue dots). The dotted vertical line represents the minimum of the curve, in correspondence to $\lambda\sim 3\cdot 10^{13}\,\text{m}\sim 200\,\text{AU}$
3.4 Comparison with theoretical estimates
As stated in the previous section, the inclusion of the 1PN acceleration in the equations of motion implies an additional assumption; namely, that any correction to the GR expression is subleading with respect to Eq. (3) and, hence, negligible in our fit. To show this, we compared our results obtained using Eq. (3) with the analytical expressions derived in the literature for some specific theories.
Alves et al. (2024) derived the 1PN acceleration in massive Brans-Dicke theory, which results in exactly the same expression as Eq. (3), as long as we identify $\alpha=a_{1}/2$, where $a_{1}=2\varphi/(2\omega_{0}+3)$, and set the reduced mass to $\eta=0$, which is clearly a good approximation for the SgrA-S22 system.
In Tan and Lu (2024) an analytical expression for the 1PN acceleration in $f(R)$ gravity is obtained and reported in their Eq. $(17)$. We used this expression to show that our upper limits on $\alpha$ are not affected by this difference in the 1PN expansion, at least when the uncertainty on $\alpha$ is the smallest (i.e., for $\lambda=3\cdot 10^{13}\,\rm m)$. In Fig. 3, the comparison between the posteriors is reported, showing that the upper limit on $\alpha$ is only changed by a factor of $2$ when the full expression for $f(R)$ gravity is used.
Tan and Lu (2024) also showed that the use of a multiple-star fit (specifically including S22, S29,29, and S5555) with the 11PN acceleration for $f(R)$ gravity derived in Eq. $(17)$ could potentially break the degeneracy between $M$ and $\alpha$ in the large $\lambda$ limit. This could produce a stringent upper limit on the fifth force intensity for $\lambda>10^{15}\,\rm m$ as well.
In addition, Losada et al. (2025) demonstrated that combining S22 motion with S6262 lensing observations can also potentially break the degeneracy between the parametrized PN parameters $\beta$ and $\gamma$. However, we leave a multi-star analysis for a future work.
Figure 3: Posterior probability density of $|\alpha|$ using the 11PN expression in Eq. (3) (black curve) versus the 11PN expansion for $f(R)$ gravity derived in Tan and Lu (2024) (red curve), when $\lambda=3\cdot 10^{13}$ m.
4 Conclusions
In this paper, we update the current constraints on the fifth force intensity at the GC using GRAVITY data for S22 from $2017$ to $2022$, including the pericenter passage. These data have allowed us to significantly improve previous estimates on the same effect, giving a $95\%$ confidence level curve that is one order of magnitude below the previous estimates (see Figure 2). Specifically, three different behaviors are found in the posterior distribution of $\alpha$, according to the value of the length scale of the new interaction $\lambda$ compared to S22 orbital range.
The minimum value of $\alpha$ is found for $\lambda=3\cdot 10^{13}\,\rm m\,(\sim 200\,AU)$ where $|\alpha|<0.003$. For comparison, in correspondence with the minimum found by Hees and others (2017), $\lambda=150\,\rm AU\sim 2.2\cdot 10^{13}\,m$, we found $|\alpha|<0.0031$. We also showed that the 11PN expansion used in this work coincides with the expression developed for massive Brans-Dicke theory and that additional terms proportional to $\alpha$ in the 11PN acceleration for $f(R)$ theories are subdominant with respect to the expression used in this work and thereby negligible, as the upper limit on $|\alpha|$ is unaffected.
A complete analysis including all S-stars, specifically those with apocenter passage farther away than S22, is left for future work, with the aim of further improving the confidence level curve. This would allow us to possibly expand the range of $\lambda$ and obtain additional and meaningful constraints at $\lambda\gtrsim 8\cdot 10^{14}\,\rm m$.
Acknowledgements.
A.F. and F.V. would like to thank Aurélien Hees and Laura Bernard for useful comments and fruitful discussions during the preparation of this work. We are very grateful to our funding agencies (MPG, ERC, CNRS [PNCG, PNGRAM], DFG, BMBF, Paris Observatory [CS, PhyFOG], Observatoire des Sciences de l’Univers de Grenoble, and the Fundação para a Ciência e a Tecnologia), to ESO and the Paranal staff, and to the many scientific and technical staff members in our institutions, who helped to make NACO, SINFONI, and GRAVITY a reality. This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 101007855. We acknowledge the financial support provided by FCT/Portugal through grants 2022.01324.PTDC, PTDC/FIS-AST/7002/2020, UIDB/00099/2020 and UIDB/04459/2020. J.S. acknowledge the National Science Foundation of China (12233001) and the National Key R&D Program of China (2022YFF0503401)
Appendix A Details of the numerical integration
The numerical integration of the equation of motion is performed using the Python library scipy.integrate.solve_ivp with a Runge-Kutta 5(4) algorithm, which means that the steps are evaluated using a fifth-order method, while the error is controlled assuming the accuracy of the fourth-order method. The convergence of the integration is ensured by looking at the conservation of energy over the entire integration period (almost two orbits in $\sim 30$ years gives $\Delta E/E\sim\mathscr{O}(10^{-10})$).
Kepler’s equation can be solved instead using a Python root finder (scipy.optimize.newton), which implements the Newton-Raphson method. The latter solves the equation with a precision of $\mathscr{O}(10^{-16})$.
Appendix B Coordinates transformations and inclusion of relativistic effects.
B.1 Coordinate transformation
The transformation from the orbital reference frame to the observer reference frame can be achieved by using the following conversion:
[…]
where $A,B,C,F,G,H$ are the Thiele-Innes parameters (Catanzarite 2010) defined as:
$\begin{split}&A=\cos\Omega\cos\omega-\sin\Omega\sin\omega\cos i,\\ &B=\sin\Omega\cos\omega+\cos\Omega\sin\omega\cos i,\\ &F=-\cos\Omega\sin\omega-\sin\Omega\cos\omega\cos i,\\ &G=-\sin\Omega\sin\omega+\cos\Omega\cos\omega\cos i,\\ &C=-\sin\omega\sin i,\\ &H=-\cos\omega\sin i\,.\end{split}$ (10)
The Cartesian coordinates $\{x_{\rm BH},y_{\rm BH},z_{\rm BH}\}$ and velocities $\{v_{x_{\rm BH}},v_{y_{\rm BH}},v_{z_{\rm BH}}\}$ are those obtained from the numerical integration. For a more detailed discussion on how the coordinate system $\{x^{\prime},y^{\prime},z_{\rm obs}\}$ and the above transformation are defined, we refer to Figure 1 and Appendix B of Grould et al. (2017).
B.2 Relativistic effects and Rømer’s delay
To produce a better fit, there are observational effects that must be included in the model.
Rømer’s delay is the difference between the time of emission of the signal $t_{\rm em}$ and the actual observational dates $t_{\rm obs}$, due to the finite speed of light. To include this delay, we used the first order Taylor’s expansion of the Røemer equation, which is expressed as:
$t_{\rm em}=t_{\rm obs}-\frac{z_{\rm obs}(t_{\rm obs})}{1+v_{z_{\rm obs}}(t_{\rm obs})}\,.$ (11)
[…]
Moreover, there are two relativistic effects that must be taken into account when S22 approaches the periastron: the relativistic Doppler shift and the gravitational redshift. Both induce a shift in the spectral lines of S22 that affects the radial velocity measurements. The former is given by
$1+z_{D}=\frac{1+v_{z_{\rm obs}}}{\sqrt{1-v^{2}}}\,$, (12)
while the gravitational redshift is defined as
$1+z_{\rm G}=\frac{1}{\sqrt{1-2U(r_{\rm em})}}\,$, (13)
where $U(r_{\rm em})$ is the potential in Eq. (1) evaluated at the time of emission $t_{\rm em}$.
The two shifts can be combined using Eq. (D.13) of Grould et al. (2017) to obtain the total radial velocity
$V_{R}\approx\frac{1}{\sqrt{1-\epsilon}}\cdot\frac{1+v_{z_{\rm obs}}/\sqrt{1-\epsilon}}{\sqrt{1-v^{2}/(1-\epsilon)}}-1\,.$ (14)
where $\epsilon=2U({r_{\rm em}})$.
In the total space velocity, $v=|\textbf{v}|$, we must also add a correction due to Solar System motion. We followed the most recent work of Reid and Brunthaler (2020) and take a proper motion of Sgr A of
$\begin{split}&v_{x}^{\rm SSM}=-5.585\,\rm mas/yr=6.415\cos(209.47^{\circ})\,mas/yr\,,\\ &v_{y}^{\rm SSM}=-3.156\,\rm mas/yr=6.415\sin(209.47^{\circ})\,mas/yr\,.\end{split}$ (15)